OpenStudy (anonymous):

Solve. 2x2 + 3x = 20

OpenStudy (smarty):

Is that 2x^2+3x=20?

OpenStudy (zepp):

Do you know how to use the quadratic formula?

OpenStudy (anonymous):

Yes smarty

OpenStudy (anonymous):

2x^2 + 3x - 20 = 0 (2x-5)(x+4)

OpenStudy (anonymous):

Use the Quadradic formula or you could try: or factor into (2x-5)(x+4)

OpenStudy (smarty):

\[-b^{2}\pm \sqrt{b ^{2}-4ac}\div2a\]

OpenStudy (anonymous):

so your zeros are at x=(5/2) and x= -4

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You don't need quadratic formula. You can just start with (2x)(x) and then fill in the blanks from there. No problem =)